An investment of Ksh1,000,000 in an account accumulated to Ksh 2,500,000 after 5 years.

(i) State the accumulation factor A(0,5) . (2 marks)

(ii) (a) Find the simple annual interest rate which would give the accumulation

factor in part (i). (3 marks)

(b) Find the annual compound interest rate which would give the

accumulation factor in part (i). (3 marks)


1
Expert's answer
2022-01-07T10:28:05-0500

i)

"A(0,5)=\frac{2500000}{1000000}=2.5"


ii)

for simple interest:

"FV=PV(1+rt)"

"1+rt=2.5"


simple annual interest rate:

"r=1.5\/t=1.5\/5=0.3"


iii)

for annual compound interest:

"FV=PV(1+r)^t"

"(1+r)^t=2.5"


annual compound interest rate:

"r=2.5^{1\/t}-1=2.5^{1\/5}-1=0.2"